OpenStudy (anonymous):

Which value for x makes the sentence true? 8x + 3 = 11 x = 1 x = 3 x = 8 x = 11

OpenStudy (happy_to_help):

try all of them out and you will find ur answer

OpenStudy (anonymous):

b

OpenStudy (happy_to_help):

wrong its obvious 8+3=11 so the answer is 1

OpenStudy (anonymous):

okay

pooja195 (pooja195):

Wrong.

OpenStudy (lizz123):

8x + 3 = 11 So you see the add sign you have to do the opposite of addition which is subtraction. In this case you have to subtract 3 both sides. 8x + 3 = 11 8x - 3 = - 3 8x ________ = ___ 8x = 8 Now, you have to divide by 8 8x = 8 ____ = ____ 8 8 So 8x over 8 is cancelled and 8 divided by 8 equals 1. Therefore, your answer will be 1. A.1

OpenStudy (anonymous):

thank you :D it was right :D

OpenStudy (lizz123):

Your very welcome. Tag me if you need any more help.

OpenStudy (abb0t):

A